# On the future cover of a sofic shift

### v1

@@ 1. Introduction
The main new result of the paper is a theorem which out of a given canonical cover produces a new one. To appreciate this, recall that Krieger proved the uniqueness of the conjugacy $\psi_{\mathbb{K}}$ in the diagram above. Without this additional requirement it is obvious that a canonical cover is not unique up to isomorphism; for example the disjoint union of two copies of the future cover will again be canonical in the sense used so far, but not with unique lifts of conjugacies. In the following we shall therefore refer to a construction of covers as weakly canonical when it gives lifts of conjugacies, but possibly lifts that are not unique, while a construction which gives rise to unique lifts of conjugacies, as the future cover does, will be called a strongly canonical cover. The main result of the paper, Theorem 4.8, describes a method to produce a strongly canonical cover out of a given weakly canonical one.
An important virtue of the construction is that it produces new strongly canonical covers of sofic shifts. Up to now only two strongly canonical covers for sofic shifts have been known; at least when we for simplicity ignore the doubling coming from the passage from right-resolving to left-resolving covers that can be achieved by ’time-reversal’. Besides the future cover it is the minimal right-resolving cover, and only for irreducible sofic shifts. It is therefore interesting to compare the strongly canonical covers that can be produced using Theorem 4.8 with the future cover and, in the irreducible case, the minimal right-resolving cover. For this we introduce a method to find the future cover of a sofic shift starting from an arbitrary given presentation. This is based on a variation of the so-called subset construction described in §3.3 of [LM]. More specifically, it is shown that a restriction to a subset of the vertexes in the subset construction, applied to an arbitrary presentation of the sofic shift, gives rise to a labeled graph whose merged graph is isomorphic to the future cover. Thus the future cover is the factor obtained from a subgraph of the subset construction by identifying vertexes with the same follower set. This fact is used to produce the examples occurring in the paper, and makes it easy to show by examples that the new strongly canonical covers that can be obtained from Theorem 4.8 in general are different both from the future cover, and in the irreducible case, also the minimal right-resolving cover.
@@ 2.1. Setup and notation
Let $G$ be a finite directed graph with edges (or arrows) $E_{G}$ and vertexes $V_{G}$. For $e\in E_{G}$ let $s_{G}(e)\in V_{G}$ be the start vertex and $t_{G}(e)$ the terminal vertex of $e$. We shall always assume that $G$ does not contain sinks or sources; i.e. that $s_{G}^{-1}(v)\neq\emptyset$ and $t_{G}^{-1}(v)\neq\emptyset$ for all $v\in V_{G}$. We extend the definition of $t_{G}$ to finite and left-infinite paths in $G$, and the definition of $s_{G}$ to finite or right-infinite paths in $G$, in the obvious way. We denote the edge shift of $G$ by $X_{G}$. Thus $X_{G}$ consists of the bi-infinite sequences
@@
Let $A$ be a finite set (the alphabet) and $L_{G}:E_{G}\to A$ a map considered as a labelling of the edges in $G$. We extend $L_{G}$ to finite or infinite paths in the natural way. In particular, when $x:=(e_{i})_{i\in\mathbb{Z}}\in X_{G}$,
@@
The first lemma deals with factor codes, right-resolving graphs and right-closing factor codes. We refer to page 18, Definition 3.3.1 and Definition 8.1.8 in [LM], respectively.
@@ 2.2. Regular vertexes and rays
The following labeled graph $(G,L_{G})$ is right-resolving and follower-separated. But it is not regular; the vertex 1 is not regular.
112233aabbaaaaddcc
@@
Every synchronized graph in the sense of Jonoska, cf. §4 in [Jo], is regular. By Lemma 3.3.15, Proposition 3.3.16 and Corollary 3.3.19 in [LM] the minimal right-resolving presentation of an irreducible sofic shift is synchronizing in the sense of Jonoska and therefore also regular. The labeled graph in Figure 3 depicts the future cover of the even shift and it is regular, but it is not synchronizing. To give an irreducible example consider the following labeled graph which is strongly connected, right-resolving, and presents the even shift.
@@
An element $x\in X_{H}$ will be called regular when $L_{H}:\mathbb{U}(x)\to\mathbb{U}(L_{H}(x))$ is surjective. Then the following clearly holds.
@@ 2.3. The future cover and Krieger’s theorem
Thus $(\mathbb{K}(Y),L_{\mathbb{K}(Y)})$ is a presentation of $Y$, and it is known as the future cover of $Y$, cf. Exercise 3.2.8 on page 75 of [LM].  [^1] Krieger introduced the word ’future’ in [Kr1], and called an equivalent version of $(\mathbb{K}(Y),L_{\mathbb{K}(Y)})$ ’the future state chain’. The terminology ”future cover” was used in [BKM], but only for a component of $(\mathbb{K}(Y),L_{\mathbb{K}(Y)})$, and only for transitive sofic shifts. In Exercise 3.2.8 of [LM] the quotation marks were deleted and $(\mathbb{K}(Y),L_{\mathbb{K}(Y)})$ was called the future cover for general sofic shifts.
@@
We must check that
$f_{\mathbb{K}(Y)}\left(t_{\mathbb{K}(Y)}(\alpha_{Y}(y)_{(-\infty,k]})\right)=F(\sigma^{k+1}(\alpha_{Y}(y)))$
@@
$f_{\mathbb{K}(Y)}\left(t_{\mathbb{K}(Y)}(\alpha_{Y}(y)_{(-\infty,k]})\right)=f_{\mathbb{K}(Y)}\left(s_{\mathbb{K}(Y)}(\alpha_{Y}(y)_{k+1})\right)$
$=f_{\mathbb{K}(Y)}(F(\sigma^{k+1}(y)))=F(\sigma^{k+1}(y))=F(\sigma^{k+1}(\alpha_{Y}(y))).$
@@
Let $v\in V_{G}$. Since $v$ is regular there is an $x\in X_{G}$ such that $t_{G}(x_{-\infty,-1]})=v$ and $f_{G}(v)=F(L_{G}(x))$. Define $\theta(v)$ such that
$\theta(v):=F(L_{G}(x))\in V_{\mathbb{K}(Y)}.$
This is well-defined: If $x^{\prime}$ has the same properties as $x$, it follows that $F(L_{F}(x^{\prime}))=f_{G}(v)=F(L_{G}(x))$.
@@
Hence $\theta(t_{G}(e))=F(L_{G}(x^{\prime}))$ and $F(L_{G}(x^{\prime}))=\frac{F(L_{G}(x))}{L_{G}(e)}$. There is therefore an edge $e^{\prime}$ in $\mathbb{K}(Y)$ from $\theta(s_{G}(e))=F(L_{G}(x))$ to $\theta(t_{G}(e))=F(L_{G}(x^{\prime}))$ with label $L_{\mathbb{K}(Y)}(e^{\prime})=L_{G}(e)$. Since $(\mathbb{K}(Y),L_{\mathbb{K}(Y)})$ is right-resolving this arrow is unique and we set $\theta(e)=e^{\prime}$. Then $\theta:G\to\mathbb{K}(Y)$ is a labeled-graph morphism.
@@ 3. Additional properties of the future cover
Let $(H,L_{H})$ be a labeled graph. A subset $U\subseteq V_{H}$ of vertexes in $H$ is a hereditary subset when $e\in E_{H},\ s_{H}(e)\in U\ \Rightarrow\ t_{H}(e)\in U$. When this holds we let $H_{U}$ denote the graph with $V_{H_{U}}:=U$ and $E_{H_{U}}:=\left\{e\in E_{H}:\ t_{H}(e)\in U\right\}$. By restriction the labeling $L_{H}$ defines a labeling of $H_{U}$ and we denote the resulting labeled graph by $(H_{U},L_{H})$. We call this a heritary labeled subgraph of $(H,L_{H})$. Lemma 2.18 can then be improved to the following.
@@
Assume $v,w\in V_{H}$ and $p_{H}(v)=p_{H}(w)$. Since $v$ is regular there is a $z\in X_{H}$ such that $t_{H}(z_{(-\infty,-1]})=v$ and $f_{H}(v)=F(L_{H}(z))$. Then $L_{H}(z_{(-\infty,-1]})\in p_{H}(v)=p_{H}(w)$, so there is a $u\in X_{H}(-\infty,-1]$ such that $t_{H}(u)=w$ and $L_{H}(u_{(-\infty,-1]})=L_{H}(z_{(-\infty,-1]})$. It follows that $f_{H}(w)\subseteq F(L_{H}(z))=f_{H}(v)$. By symmetry $f_{H}(v)\subseteq f_{H}(w)$ and hence $f_{H}(v)=f_{H}(w)$, which implies that $v=w$ since we assume that $(H,L_{H})$ is follower-separated. ∎
@@ 3.3. The relation between the future cover and the follower set graph
The proof uses the observation that two words $w_{1},w_{2}\in\mathbb{W}(Y)$ have the same follower sets as defined in [LM] if and only if $F(w_{1})=F(w_{2})$. Let $y\in Y$. By Lemma 2.20 there is an $N\in\mathbb{N}$ such that $F(y)=F(y_{[-n,-1]})$ for all $n\geq N$. If $z\in Y$ and $F(y)=F(z)$ there is an $M\in\mathbb{N}$ such that $F(z_{[-n,-1]})=F(y_{[-n,-1]})$ for all $n\geq M$. It follows that we can define a map $\chi:V_{\mathbb{K}(Y)}\to V_{G}$ such that $\chi(F(y))=F(y_{[-n,-1]})$ for all large $n\in\mathbb{N}$. This map, $\chi$, is clearly injective. For $y,z\in Y$ there is a labeled arrow $F(y)\overset{a}{\to}F(z)$ in $(\mathbb{K}(Y),L_{\mathbb{K}(Y)})$ if and only if $F(z)=\frac{F(y)}{a}$ if and only if $F(z_{[-n,-1]})=\frac{F(y_{[-n,-1]})}{a}$ for all large $n$ if and only of there is a labeled arrow $\chi(F(y))\overset{a}{\to}\chi(F(z))$ in $G$. This shows that $\chi$ is an embedding of labeled graphs. To show that $\chi(V_{\mathbb{K}(Y)})$ is a hereditary subset, let $w\in\mathbb{W}(Y)$ and assume that $\chi(F(y))\overset{a}{\to}F(w)$ in $G$. Then $F(w)=\frac{F(y)}{a}$, and hence $F(w)=F(z)$, where $z\in Y$ is any element with $z_{(-\infty,-1]}=y_{(-\infty,-1]}a$. It follows that $F(w)=\chi(F(z))$.
Since $(\mathbb{K}(Y),L_{\mathbb{K}(Y)})$ is regular by Lemma 2.13 and since the range of $\chi$ is a hereditary subgraph of $(G,\mathcal{L})$ it follows that $\chi(V_{\mathbb{K}(Y)})$ consist only of regular vertexes in $G$. Conversely, if $w\in\mathbb{W}(Y)$ and $F(w)$ is a regular vertex in $G$, it follows from the defining relation (1) that $F(w)=F(z)$ for some $z\in Y$, proving that $F(w)=\chi(F(z))\in\chi(V_{\mathbb{K}(Y)})$.
@@ 4.2. Subset constructions for sofic shifts
To construct an inverse map $\kappa^{-1}:X_{H^{\prime\prime}}\to X_{H}^{L_{H}}$ to $\kappa$, let $x=(x_{i})_{i\in\mathbb{Z}}\in X_{H^{\prime\prime}}$. Since $(H,L_{H})$ is right-resolving, $s_{H^{\prime\prime}}(x_{i})\geq\#s_{H^{\prime\prime}}(x_{i+1})$, and we choose $N\in\mathbb{N}$ such that $s_{H^{\prime\prime}}(x_{i})=\#s_{H^{\prime\prime}}(x_{N})$ for all $i\leq N$. For each $v\in s_{H^{\prime\prime}}(x_{N})$ there is a unique path $\gamma^{v}\in X_{H}$ such that $s_{H}(\gamma^{v}_{N})=v$ and $L_{H}(\gamma_{v})=L_{H^{\prime\prime}}(x)$. Setting
$\kappa^{-1}(x):=\left\{\gamma_{v}:\ v\in s_{H^{\prime\prime}}(x_{N})\right\}$
@@
$L_{H^{\prime\prime}}\circ\beta_{Y}=\operatorname{id}_{Y}.$
From the proof of Lemma 4.2 we get the following.
@@
Let $y\in Y$. Then $\beta_{H}(y)\in X_{H^{\prime\prime}}$ is the unique element with the properties that $L_{H^{\prime\prime}}(\beta_{H}(y))=y$ and $s_{H^{\prime\prime}}(\beta_{H}(y)_{j})=\left\{s_{H}(x_{j}):\ x\in L_{H}^{-1}(y)\right\}$ for all $j\in\mathbb{Z}$.
Lemma 4.5.
Let $C$ be a component in $H^{\prime\prime}$ such that there is an element $y\in Y$ with $\beta_{H}(y)$ backward asymptotic to $X_{C}$, and let $z\in X_{H^{\prime\prime}}$ be an element backward transitive to $X_{C}$. Then $z=\beta_{H}(L_{H^{\prime\prime}}(z))$.
@@
Set $y^{\prime}:=L_{H^{\prime\prime}}(z)$. Then $y^{\prime}$ is backward transitive to $L_{H^{\prime\prime}}(X_{C})$. Let $C^{\prime}$ be the component in $H^{\prime\prime}$ with the property that $\beta_{H}(y^{\prime})$ is backward asymptotic to $X_{C^{\prime}}$. Note that the number
$M(C):=\#s_{H^{\prime\prime}}(x_{j})$
is the same for all elements $x\in X_{C}$ and all $j\in\mathbb{Z}$. Since $L_{H^{\prime\prime}}(z)=y^{\prime}$ we have that
(11) $s_{H^{\prime\prime}}(z_{j})\subseteq s_{H^{\prime\prime}}(\beta_{H}(y^{\prime})_{j})$
for all $j\in\mathbb{Z}$, implying that $M(C)\leq M(C^{\prime})$ since $z$ is backward asymptotic to $X_{C}$ and $\beta_{H}(y^{\prime})$ to $X_{C^{\prime}}$. On the other hand, if $w$ is a word in $L_{H^{\prime\prime}}(X_{C})$ it occurs infinitely often to the left in $y^{\prime}$. Since $L_{H^{\prime\prime}}^{-1}(w)$ is a finite set and $L_{H^{\prime\prime}}(\beta_{H}(y^{\prime}))=y^{\prime}$ it follows that $w=L_{H^{\prime\prime}}(w^{\prime})$ for some word $w^{\prime}$ appearing infinitely often to the left in $\beta_{H}(y^{\prime})$. Then $w^{\prime}\in\mathbb{W}(X_{C^{\prime}})$ and we conclude therefore that $L_{H^{\prime\prime}}(X_{C})\subseteq L_{H^{\prime\prime}}(X_{C^{\prime}})$. It follows that there is a $z^{\prime}\in X_{C^{\prime}}$ such that $y$ is backward asymptotic to $L_{H^{\prime\prime}}(z^{\prime})$, and there is therefore a $N\in\mathbb{Z}$ such that $s_{H^{\prime\prime}}(z^{\prime}_{j})\subseteq s_{H^{\prime\prime}}(\beta_{H}(y)_{j})$ for all $j\leq N$. It follows that $M(C^{\prime})\leq M(C)$ and therefore also that $M(C)=M(C^{\prime})$. Then (11) implies that $z=\beta_{H}(y^{\prime})=\beta_{H}(L_{H^{\prime\prime}}(z))$. ∎
@@
To simplify the notation, set $L_{H^{\prime}}:=L_{H^{\prime\prime}}|_{H^{\prime}}$.
Lemma 4.7.
$X_{H^{\prime}}=\overline{\beta_{H}(Y)}$. In particular, $(X_{H^{\prime}},L_{H^{\prime}})$ is a right-resolving presentation of $Y$.
Proof.
By definition of $H^{\prime}$ every element of $X_{H^{\prime}}$ can be approximated by elements $z$ with the property that $z$ is backward transitive to $X_{C}$, where $C$ is the component in $H^{\prime\prime}$ with the property that $\beta_{H}(y)$ is backward asymptotic to $X_{C}$ for some $y\in Y$. Since $z\in\beta_{H}(Y)$ by Lemma 4.5, it follows that $X_{H^{\prime}}\subseteq\overline{\beta_{H}(Y)}$. On the other hand, every element of $\beta_{H}(Y)$ is backward asymptotic to some $X_{C}$, where $C$, by definition, is a component in $H^{\prime}$. Hence $\beta_{H}(Y)\subseteq X_{H^{\prime}}$. ∎

### v2

@@ 1. Introduction
The main new result of the paper is a theorem which out of a given canonical cover produces a new one. To appreciate this, recall that Krieger proved the uniqueness of the conjugacy $\psi_{\mathbb{K}}$ in the diagram above. Without this additional requirement it is obvious that a canonical cover is not unique up to isomorphism; for example the disjoint union of two copies of the future cover will again be canonical in the sense used so far, but not with unique lifts of conjugacies. In the following we shall therefore refer to a construction of covers as weakly canonical when it gives lifts of conjugacies, but possibly lifts that are not unique, while a construction which gives rise to unique lifts of conjugacies, as the future cover does, will be called a strongly canonical cover. The main new result of the paper, Theorem 4.6, describes a method to produce a strongly canonical cover out of a given weakly canonical one.
An important point concerning the construction is that it produces new strongly canonical covers of sofic shifts. Up to now only two strongly canonical covers for sofic shifts have been known; at least when we for simplicity ignore the doubling coming from the passage from right-resolving to left-resolving covers that can be achieved by ’time-reversal’. Besides the future cover it is the minimal right-resolving cover, and only for irreducible sofic shifts. It is therefore interesting to compare the strongly canonical covers that can be produced using Theorem 4.6 with the future cover and, in the irreducible case, the minimal right-resolving cover. For this we introduce a method to find the future cover of a sofic shift starting from an arbitrary given presentation. This is based on a variation of the so-called subset construction described in §3.3 of [LM]. More specifically, it is shown that a restriction to a subset of the vertexes in the subset construction, applied to an arbitrary presentation of the sofic shift, gives rise to a labeled graph whose merged graph is isomorphic to the future cover. Thus the future cover is the factor obtained from a subgraph of the subset construction by identifying vertexes with the same follower set. This fact is used to produce the examples occurring in the paper, and makes it easy to show by examples that the new strongly canonical covers that can be obtained from Theorem 4.6 in general are different both from the future cover, and in the irreducible case, also the minimal right-resolving cover.
@@ 2.1. Setup and notation
Let $G$ be a finite directed graph with edges (or arrows) $E_{G}$ and vertexes $V_{G}$. For $e\in E_{G}$ let $s_{G}(e)\in V_{G}$ be the start vertex and $t_{G}(e)$ the terminal vertex of $e$. We shall always assume that $G$ does not contain sinks or sources; i.e. that $s_{G}^{-1}(v)\neq\emptyset$ and $t_{G}^{-1}(v)\neq\emptyset$ for all $v\in V_{G}$. We extend the definition of $t_{G}$ to finite and left-infinite paths in $G$, and the definition of $s_{G}$ to finite and right-infinite paths in $G$, in the obvious way. We denote the edge shift of $G$ by $X_{G}$. Thus $X_{G}$ consists of the bi-infinite sequences
@@
Let $A$ be a finite set (the alphabet) and $L_{G}:E_{G}\to A$ a map considered as a labelling of the edges in $G$. We extend $L_{G}$ to finite or infinite paths in the natural way. In particular, when $x=(e_{i})_{i\in\mathbb{Z}}\in X_{G}$,
@@
The first lemma deals with factor codes, right-resolving graphs and right-closing factor codes. We refer to page 18, Definition 3.3.1, and Definition 8.1.8 in [LM], respectively.
@@ 2.2. Regular vertexes and rays
The following labeled graph $(G,L_{G})$ is right-resolving and follower-separated. But it is not regular; the vertex a is not regular.
aabbcc112211114433
@@
Every synchronizing graph in the sense of Jonoska, cf. §4 in [Jo], is regular. By Lemma 3.3.15, Proposition 3.3.16 and Corollary 3.3.19 in [LM] the minimal right-resolving presentation of an irreducible sofic shift is synchronizing in the sense of Jonoska and therefore also regular. The labeled graph in Figure 3 depicts the future cover of the even shift and it is regular, but it is not synchronizing. To give an irreducible example consider the following labeled graph which is strongly connected, right-resolving, and presents the even shift.
@@
An element $x\in X_{H}$ will be called regular when $L_{H}:\mathbb{U}(x)\to\mathbb{U}(L_{H}(x))$ is surjective. By Lemma 2.5 the following clearly holds.
@@ 2.3. The future cover and Krieger’s theorem
Thus $(\mathbb{K}(Y),L_{\mathbb{K}(Y)})$ is a presentation of $Y$, and it is known as the future cover of $Y$, cf. Exercise 3.2.8 on page 75 of [LM].  [^1] Krieger introduced the word ’future’ in [Kr1], and called an equivalent version of $(\mathbb{K}(Y),L_{\mathbb{K}(Y)})$ ’the future state chain’. The terminology ”future cover” was used in [BKM], but only for a component of $(\mathbb{K}(Y),L_{\mathbb{K}(Y)})$, and only for transitive sofic shifts. In Exercise 3.2.8 of [LM] the quotation marks were deleted and $(\mathbb{K}(Y),L_{\mathbb{K}(Y)})$ was called the future cover for general sofic shifts. In some papers on sofic shifts the future cover is called the Krieger cover and the minimal right-resolving cover is called the Fischer cover.
@@
Since $L_{\mathbb{K}(Y)}(\alpha_{Y}(y))=y$ we must check that
$f_{\mathbb{K}(Y)}\left(t_{\mathbb{K}(Y)}(\alpha_{Y}(y)_{(-\infty,k]})\right)=F(\sigma^{k+1}(y))$
@@
$f_{\mathbb{K}(Y)}\left(t_{\mathbb{K}(Y)}(\alpha_{Y}(y)_{(-\infty,k]})\right)=f_{\mathbb{K}(Y)}\left(s_{\mathbb{K}(Y)}(\alpha_{Y}(y)_{k+1})\right)=f_{\mathbb{K}(Y)}(F(\sigma^{k+1}(y)))=F(\sigma^{k+1}(y)).$
@@
Let $v\in V_{G}$. Since $v$ is regular there is an $x\in X_{G}$ such that $t_{G}(x_{-\infty,-1]})=v$ and $f_{G}(v)=F(L_{G}(x))$. Define $\theta(v)\in V_{\mathbb{K}(Y)}$ such that
$\theta(v):=F(L_{G}(x)).$
This is well-defined: If $x^{\prime}$ has the same properties as $x$, it follows that $F(L_{G}(x^{\prime}))=f_{G}(v)=F(L_{G}(x))$.
@@
Hence $\theta(t_{G}(e))=F(L_{G}(x^{\prime}))$ and $F(L_{G}(x^{\prime}))=\frac{F(L_{G}(x))}{L_{G}(e)}$. There is therefore an edge $e^{\prime}$ in $\mathbb{K}(Y)$ from $\theta(s_{G}(e))=F(L_{G}(x))$ to $\theta(t_{G}(e))=F(L_{G}(x^{\prime}))$ with label $L_{\mathbb{K}(Y)}(e^{\prime})=L_{G}(e)$. Since $(\mathbb{K}(Y),L_{\mathbb{K}(Y)})$ is right-resolving this arrow is unique and we set $\theta(e):=e^{\prime}$. Then $\theta:G\to\mathbb{K}(Y)$ is a labeled-graph morphism.
@@ 3. Additional properties of the future cover
Let $(H,L_{H})$ be a labeled graph. A subset $U\subseteq V_{H}$ of vertexes in $H$ is a hereditary subset when $e\in E_{H},\ s_{H}(e)\in U\ \Rightarrow\ s_{H}(e)\in U$. When this holds we let $H_{U}$ denote the graph with $V_{H_{U}}:=U$ and $E_{H_{U}}:=\left\{e\in E_{H}:\ t_{H}(e)\in U\right\}$. By restriction the labeling $L_{H}$ defines a labeling of $H_{U}$ and we denote the resulting labeled graph by $(H_{U},L_{H})$. We call this a heritary labeled subgraph of $(H,L_{H})$. Lemma 2.18 can then be improved to the following.
@@
Assume $v,w\in V_{H}$ and $p_{H}(v)=p_{H}(w)$. Since $v$ is regular there is a $z\in X_{H}$ such that $t_{H}(z_{(-\infty,-1]})=v$ and $f_{H}(v)=F(L_{H}(z))$. Then $L_{H}(z_{(-\infty,-1]})\in p_{H}(v)=p_{H}(w)$, so there is a $u\in X_{H}(-\infty,-1]$ such that $t_{H}(u)=w$ and $L_{H}(u_{(-\infty,-1]})=L_{H}(z_{(-\infty,-1]})$. It follows that $f_{H}(w)\subseteq F(L_{H}(u))=F(L_{H}(z))=f_{H}(v)$. By symmetry $f_{H}(v)\subseteq f_{H}(w)$ and hence $f_{H}(v)=f_{H}(w)$, which implies that $v=w$ since we assume that $(H,L_{H})$ is follower-separated. ∎
@@ 3.3. The relation between the future cover and the follower set graph
The proof uses the observation that two words $w_{1},w_{2}\in\mathbb{W}(Y)$ have the same follower sets as defined in [LM] if and only if $F(w_{1})=F(w_{2})$. Let $y\in Y$. By Lemma 2.20 there is an $N\in\mathbb{N}$ such that $F(y)=F(y_{[-n,-1]})$ for all $n\geq N$. If $z\in Y$ and $F(y)=F(z)$ there is an $M\in\mathbb{N}$ such that $F(z_{[-n,-1]})=F(y_{[-n,-1]})$ for all $n\geq M$. It follows that we can define a map $\chi:V_{\mathbb{K}(Y)}\to V_{G}$ such that $\chi(F(y))=F(y_{[-n,-1]})$ for all large $n\in\mathbb{N}$. This map, $\chi$, is clearly injective. For $y,z\in Y$ there is a labeled arrow $F(y)\overset{a}{\to}F(z)$ in $(\mathbb{K}(Y),L_{\mathbb{K}(Y)})$ if and only if $F(z)=\frac{F(y)}{a}$ if and only if $F(z_{[-n,-1]})=\frac{F(y_{[-n,-1]})}{a}$ for all large $n$ if and only if there is a labeled arrow $\chi(F(y))\overset{a}{\to}\chi(F(z))$ in $G$. This shows that $\chi$ is an embedding of labeled graphs. To show that $\chi(V_{\mathbb{K}(Y)})$ is a hereditary subset, let $w\in\mathbb{W}(Y)$ and assume that $\chi(F(y))\overset{a}{\to}F(w)$ in $G$. Then $F(w)=\frac{F(y)}{a}$, and hence $F(w)=F(z)$, where $z\in Y$ is any element with $z_{(-\infty,-1]}=y_{(-\infty,-1]}a$. It follows that $F(w)=\chi(F(z))$.
Since $(\mathbb{K}(Y),L_{\mathbb{K}(Y)})$ is regular by Lemma 2.13 and since the range of $\chi$ is a hereditary subgraph of $(G,\mathcal{L})$ it follows from Lemma 2.2 that $\chi(V_{\mathbb{K}(Y)})$ consist only of regular vertexes in $G$. Conversely, if $w\in\mathbb{W}(Y)$ and $F(w)$ is a regular vertex in $G$, it follows from the defining relation (1) that $F(w)=F(z)$ for some $z\in Y$, proving that $F(w)=\chi(F(z))\in\chi(V_{\mathbb{K}(Y)})$.
@@ 4.2. Subset constructions for sofic shifts
To construct an inverse map $\kappa^{-1}:X_{H^{\prime\prime}}\to X_{H}^{L_{H}}$ to $\kappa$, let $x=(x_{i})_{i\in\mathbb{Z}}\in X_{H^{\prime\prime}}$. Since $(H,L_{H})$ is right-resolving, $s_{H^{\prime\prime}}(x_{i})\geq\#s_{H^{\prime\prime}}(x_{i+1})$, and we choose $N\in\mathbb{N}$ such that $s_{H^{\prime\prime}}(x_{i})=\#s_{H^{\prime\prime}}(x_{N})$ for all $i\leq N$. For each $v\in s_{H^{\prime\prime}}(x_{N})$ there is a unique path $\gamma^{v}\in X_{H}$ such that $s_{H}(\gamma^{v}_{N})=v$, $s_{H}(\gamma^{v}_{i})\in s_{H^{\prime\prime}}(x_{i})$ for all $i\in\mathbb{Z}$ and $L_{H}(\gamma^{v})=L_{H^{\prime\prime}}(x)$. Setting
$\kappa^{-1}(x):=\left\{\gamma^{v}:\ v\in s_{H^{\prime\prime}}(x_{N})\right\}$
@@
$L_{H^{\prime\prime}}\circ\beta_{H}=\operatorname{id}_{Y}.$
Concerning the components of $H^{\prime\prime}$ we will say that a component $C^{\prime}$ in $H^{\prime\prime}$ connects to another component $C$ when there is a path $\mu$ in $H^{\prime\prime}$ such that $s_{H^{\prime\prime}}(\mu)\in C^{\prime}$ and $t_{H^{\prime\prime}}(\mu)\in C$. Note that for any component $C$ in $H^{\prime\prime}$ the cardinality #F\#F of the elements $F$ of $C\subseteq 2^{V_{H}}$ is the same; this number is the multiplicity of $C$ and we denote it by $M(C)$.
@@
Let $C$ be a component in $H^{\prime\prime}$ such that $\beta_{H}(y)$ is backward asymptotic to $X_{C}$ for some $y\in Y$. Then $X_{C}\subseteq\overline{\beta_{H}(Y)}$.
@@
Let $y\in Y$ and let $C$ be the component in $H^{\prime\prime}$ for which there is a $N\in\mathbb{Z}$ such that $\beta_{H}(y)_{j}\in C$ for all $j\leq N$. It follows that there is an $M<N$ such that if $z\in X_{H^{\prime\prime}}$ and $L_{H^{\prime\prime}}(z)_{[M,\infty)}=y_{[M,\infty)}$, then $s_{H^{\prime\prime}}(z_{N})\subseteq s_{H^{\prime\prime}}(\beta_{H}(y)_{N})$. (If not, a compactness argument would produce an element $z^{\prime}\in X_{H^{\prime\prime}}$ such that $L_{H^{\prime\prime}}(z^{\prime})=y$ and $s_{H^{\prime\prime}}(z^{\prime}_{N})>\#s_{H^{\prime\prime}}(\beta_{H}(y)_{N})$. A contradiction.) Let $C^{\prime}$ be a component which connects to $C$ and has the largest multiplicity among the components that connect to $C$. Let $\gamma$ be a finite path in $H^{\prime\prime}_{C}$. We can then choose an element $z^{\prime}\in X_{H^{\prime\prime}}$ with the following properties:
  • i)
    $z^{\prime}$ is backward asymptotic to $C^{\prime}$,
  • ii)
    there is a $M^{\prime}<M$ such that $z^{\prime}_{[M^{\prime},M]}$ is a path in $H^{\prime\prime}_{C}$ with $\gamma\subseteq z^{\prime}_{[M^{\prime},M]}$, and
  • iii)
    $z^{\prime}_{[M,\infty)}=\beta_{H}(y)_{[M,\infty)}$.
We claim that
(11) $z^{\prime}=\beta_{H}(L_{H^{\prime\prime}}(z^{\prime})).$
To see this, let $C^{\prime\prime}$ be the component with the property that $\beta_{H}(L_{H^{\prime\prime}}(z^{\prime}))$ is backward asymptotic to $X_{C^{\prime\prime}}$. Then
(12) $\beta_{H}(L_{H^{\prime\prime}}(z^{\prime}))_{N}=\beta_{H}(y)_{N}.$
Indeed, $L_{H^{\prime\prime}}(\beta_{H}(L_{H^{\prime\prime}}(z^{\prime}))_{[M,\infty)}=L_{H^{\prime\prime}}(z^{\prime})_{[M,\infty)}=y_{[M,\infty)}$ by iii), which implies that
$s_{H^{\prime\prime}}(\beta_{H}(L_{H^{\prime\prime}}(z^{\prime}))_{N})\subseteq s_{H^{\prime\prime}}(\beta_{H}(y)_{N})$
by the choice of $M$, while the opposite inclusion follows from
$s_{H^{\prime\prime}}(\beta_{H}(y)_{N})=s_{H^{\prime\prime}}(z^{\prime}_{N})\subseteq s_{H^{\prime\prime}}(\beta_{H}(L_{H^{\prime\prime}}(z^{\prime}))_{N}).$
Since
(13) $s_{H^{\prime\prime}}(z^{\prime}_{j})\subseteq s_{H^{\prime\prime}}(\beta_{H}(L_{H^{\prime\prime}}(z^{\prime}))_{j})$
for all $j$ we have that $M(C^{\prime})\leq M(C^{\prime\prime})$. But it follows from (12) that $C^{\prime\prime}$ connects to $C$ and hence $M(C^{\prime\prime})\leq M(C^{\prime})$ by definition of $C^{\prime}$ and we find therefore that $M(C^{\prime\prime})=M(C^{\prime})$. It follows from this identity and (13) that $s_{H^{\prime\prime}}(z^{\prime}_{j})=s_{H^{\prime\prime}}(\beta_{H}(L_{H^{\prime\prime}}(z^{\prime}))_{j})$ for all $j$ ’close to -$\infty$’; a conclusion which implies the claim; (11). The inclusion $X_{C}\subseteq\overline{\beta_{H}(Y)}$ follows now because the path $\gamma$ is arbitrary and $\beta_{H}(Y)$ is $\sigma$-invariant. ∎
@@
To simplify notation, set $L_{H^{\prime}}:=L_{H^{\prime\prime}}|_{H^{\prime}}$.